Open Source Writers Group: Mission Statement

1. Purpose of this Document

The Open Source Writers Group Mission Statement has been written to clarify the purpose, goals, and future of the OSWG. This document exists to help other projects understand what the OSWG is about, and to help current and future OSWG members understand what we’re trying to do.

This Mission Statement will be an evolving document, changing over time as the needs of the community change and the OSWG, in turn, changes to fill those needs. This Mission Statement is not "written in stone", and the OSWG maintainers reserve the right to modify the Mission Statement when circumstances require.

Non-modifiable section: There is one non-modifiable section of the Mission Statement – the first paragraph of the Mission Summary and Related Activities. It is of utmost importance that the OSWG always remain a non-commercial and non-profit organization with a primary goal of improving the quality and quantity of free open-source and open-content documentation. This is non-negotiable.

2. Mission Summary and Related Activities

The Open Source Writers Group is a non-commercial, non-profit organization whose primary goal is to improve the overall quality and quantity of free open-source and open-content documentation. All OSWG activities and projects are related to this primary goal.

The following is a list of ways in which the OSWG works to fulfill this goal.

Open Source Writers Group: Mission Statement

2.1. Improving Project-Volunteer Communication

One method by which the OSWG works to improve open-source/open-content documentation is by helping to improve the communication between volunteer writers and editors and the open-source projects that need them. These efforts are largely related to helping projects recruit volunteers to work on project documentation. This includes (but is not limited to) a volunteer/project "registration" program through which volunteers make it known that they are willing and able to help with documentation for open-source projects. Similarly, open-source projects can also register to make it known that they are looking for people to help with documentation. The OSWG thus provides a central point of contact for projects and volunteers, helping to match each with the other. Currently the "matching" process is a fairly passive role, but the OSWG will take a more active role when time and circumstance allow.

2.2. Improving Project-Project Communication

The OSWG also provides a forum in which people working on various open-source documentation projects can share ideas, experience, and knowledge. Enabling this sort of knowledge exchange is of key importance – documentation projects often encounter many of the same or similar problems and challenges. Being able to discuss these problems with others who have already worked through those challenges can be very valuable.

Fostering increased cooperation and collaboration between disparate documentation projects is one of the main methods through which the OSWG hopes to improve open-source/open-content documentation. The sharing of ideas, experience, and knowledge is a strong benefit to each project, and to the community as a whole.

2.3. Providing Free Collaborative Documentation Tools and Repository

The OSWG also works to provide useful and open collaborative documentation tools and a centralized documentation repository for any free open-source or open-content documentation. These tools are available for any writer or group of writers who wish to use them, so long as the documentation license falls within the acceptable range of licenses as outlined in the OSWG Licensing Policy.

2.4. Working to Improve the Usability of Existing Free Documentation

As many people who are experienced with open-source know, there is a wealth of information available both online and off. One problem most often encountered by newcomers to the open-source, however, is that it can be very difficult to find the specific information required to solve a particular problem. It’s the "needle in a haystack" issue, more recently called "the Information Fire Hose" (a term coined by Alan Cox). There is so much information available, that sifting through it all can be both frustrating and time-consuming.

The OSWG hopes to improve this situation by helping the various projects that are working towards solving these and similar problems. Two such projects are the Linux Knowledge Base (linuxkb.org) and the Open Source Research Team’s Metadata project (see metalab.unc.edu/osrt).

2.5. Promoting Open-Source and Open-Content Documentation Projects

There is truth in the adage that there is "strength in numbers". The OSWG, as a "neutral" organization in the open-source/open-content community, is unaligned with any particular project, operating system, or corporation. This neutrality allows the OSWG to help promote any of the numerous documentation projects that are part of this community. The more projects the OSWG helps promote, the better it is for all involved, as the OSWG will become known as a centralized location where people can get news, information, and help with open-source and open-content documentation.

The OSWG is uniquely situated in this because the group is not "about Linux" or "about GNOME" or "about FreeBSD". The OSWG is about all of these projects, and more. At its foundation, the OSWG is simply about improving all open-source and open-content documentation, and can therefore work to help all open-source projects, rather than just a specific subset of them.

2.6. Providing an Open-Source Documentation News and Announcements Resource

As community awareness of the OSWG grows, the OSWG website will also grow to become a central news and information site for open-source and open-content documentation. The site maintainer(s) will accept news pointers, announcements, press releases, and the like, from contributors and post these to the site in as timely a fashion as possible.

Again, because of the "neutral" nature of the OSWG, posted news and announcements can be about any open-source or open-content documentation project, not just those pertaining to a particular subset of the open-source community.

3. Getting Involved

If you are interested in getting involved with the OSWG, please visit our website at http://www.oswg.org. Also feel free to contact any of the current OSWG maintainers at the following email addresses:

<editor@oswg.org>

Note: general coordinator/contact <webmaster@oswg.org>

Note: website maintainer <cvs-maintainer@oswg.org>

Note: CVS system maintainer <list-maintainer@oswg.org>

Note: mailing list maintainer

It is also recommended that you subscribe to the main OSWG discussion list, oswg-discuss. To subscribe, send mail with subscribe in the subject line to

<oswg-discuss-request@oswg.org>

List archives can be found through the main OSWG website at http://www.oswg.org. If you have any questions, comments, or suggestions, please contact the current OSWG general maintainer at:

<editor@oswg.org>

.